how do you solve the problem 5x=4x+8?
Answer by nyc_function(2741)

5x = 4x+8

5x - 4x = 8

x = 8

The answer is 8.

THE PROVE:

Let x = 8

5x = 4x+8

5(8) = 4(8) + 8

40 = 32 + 8

40 = 40...It checks!!!

This confirms that the right answer is x = 8.